13.) What IQ interval captures the middle 25% of the population (n-1)?
Answer #1

Solution:

Given that,

P(-z < Z < z) = 25%

P(-z < Z < z) = 0.25

P(Z < z) - P(Z < z) = 0.25

2P(Z < z) - 1 = 0.25

2P(Z < z) = 1 + 0.25

2P(Z < z) = 1.25

P(Z < z) = 1.25 / 2 = 0.625

P(Z < 0.3186) = 0.625

z = 0.32

z = - 0.32